# Phonon-Coupled Hole-Spin Qubits in High-Purity Germanium: Design and Modeling of a Scalable Architecture

### v2

@@ Abstract
We present a design and modeling of a scalable quantum processor architecture utilizing hole-spin qubits defined in gate-controlled germanium (Ge) quantum dots, where coherent spin-phonon coupling is predicted to facilitate qubit manipulation and long-range interactions. The architecture exploits the strong, electrically tunable spin-orbit interactions intrinsic to hole states in Ge, integrated with high-quality phononic crystal cavities (PnCCs) to enable fully electrical qubit control and phonon-mediated coupling. Employing a streamlined simulation framework built upon multiband $\mathbf{k}\cdot\mathbf{p}$ modeling and finite-element methods, we quantify key performance metrics, including electrically tunable $g$-factors ranging from $1.3$ to $2.0$, spin-phonon coupling strengths up to $6.3\,\mathrm{MHz}$, phononic cavity quality factors exceeding $10^{4}$, and phonon-mediated spin relaxation times ($T_{1}$) reaching milliseconds. The proposed architecture concurrently achieves extended spin coherence and rapid gate operations through strategic electric field modulation and engineered phononic bandgap environments. Furthermore, isotopically enriched, high-purity Ge crystals—grown in-house at the University of South Dakota—significantly enhance device coherence by minimizing disorder and hyperfine interactions. This integrated approach, merging advanced materials engineering, precise spin-orbit coupling, and phononic cavity design, establishes a promising CMOS-compatible pathway toward scalable, high-fidelity quantum computing.
@@ I Introduction
Figure 1: Comparison of representative qubit technologies based on coherence time and gate speed, reflecting current experimental capabilities [1, 2, 8, 9]. Ge-based hole-spin qubits with phonon coupling are positioned in a favorable regime that combines fast, all-electrical control with long coherence, underscoring their promise for high-fidelity quantum information processing.
@@
A critical factor influencing the performance of Ge-based hole-spin qubits is material quality, particularly impurity scattering and charge noise, which significantly degrade qubit coherence. To address these challenges, this study leverages in-house capabilities at the University of South Dakota (USD) for growing detector-grade, high-purity Ge crystals, achieving impurity concentrations as low as ∼$\sim 10^{10}\,\text{cm}^{-3}$ [13, 14, 15, 16, 17, 18, 19]. This vertically integrated crystal growth infrastructure provides a robust pathway for fabricating low-disorder quantum dots, thus enabling detailed exploration of phonon-assisted spin manipulation within exceptionally clean Ge-based systems.
A distinctive strength of our proposed architecture lies in the strategic integration of isotopically enriched, high-purity 74Ge substrates, the strong and electrically tunable spin-orbit coupling inherent to hole states, and precisely engineered phononic crystals (PnCs) with embedded PnCCs. Together, these components form a robust platform for spin-phonon-based quantum information processing. The PnC structure is designed to surround the quantum dot, creating a phonon bandgap that suppresses unwanted phonon modes and isolates a target frequency range (e.g., around 6 GHz). Localized defects introduced into the PnC form PnCCs that support discrete acoustic modes resonant with the desired frequency. These modes are spatially aligned with the wavefunction of the hole-spin qubit in the Ge quantum dot, enabling coherent spin-phonon interaction via the deformation potential. While the PnC acts as an acoustic shield to minimize phonon leakage and external noise, the PnCC enhances the local phonon density of states, thereby increasing the coupling strength. This integrated approach facilitates strong, localized spin-phonon coupling essential for coherent qubit control, high-fidelity readout, and phonon-mediated qubit entanglement.
@@ II Spin-Phonon Coupling in Germanium Quantum Dots
Figure 2: Schematic showing spin-phonon coupling in a Ge quantum dot. Under an applied magnetic field along [001], spin states split by the Zeeman energy $\Delta E$, and phonons with matching energy $E_{ph}=\hbar\omega$ enable spin-flip transitions. These transitions are mediated via spin-orbit coupling and enhanced by heavy-hole/light-hole mixing in low-symmetry environments. Here, $\lvert v\rangle$ and $\lvert w\rangle$ denote the spin-down and spin-up eigenstates of the spin qubit under an external magnetic field, corresponding to the spin-split states in the Ge quantum dot.
@@
At cryogenic temperatures ($<1~{}\text{K}$) relevant for quantum computing, acoustic phonons dominate over optical modes. The primary spin relaxation mechanism in this regime is acoustic deformation potential coupling, whereby phonons modulate the local band structure and mediate transitions between spin states [7, 9]. This spin-phonon interaction plays a dual role: while it enables efficient spin initialization and readout via phonon emission or absorption, it can also introduce decoherence through relaxation ($T_{1}$) and dephasing ($T_{2}^{*}$) pathways.
To mitigate these effects and simultaneously enable coherent control, it is advantageous to operate the qubit at frequencies compatible with standard microwave electronics and engineered phononic environments. In particular, targeting the 2–6 GHz frequency range allows for integration with superconducting resonators and phononic cavities designed for enhanced spin-phonon coupling. To achieve spin-flip transitions in this range, the Zeeman splitting must fall within approximately 8 to 25 $\mu$eV.
As illustrated in Figure 3, this condition can be met by tuning the hole $g$-factor and applied magnetic field. For example, a hole $g$-factor of 1.3—realistic for strained Ge quantum dots—combined with a magnetic field of 0.25 T yields a Zeeman energy of roughly 19 $\mu$eV, placing it well within the desired operational window.
Figure 3: Zeeman energy as a function of magnetic field $B$ for various hole $g$-factors, overlaid with phonon and resonator mode energies. Each curve shows the Zeeman splitting $\Delta_{E}=g\mu_{B}B$ for hole spin qubits with tunable $g$-factors in the range $g=0.5$ to $2.0$. The shaded region indicates the usable microwave control band (2–6 GHz), corresponding to Zeeman energies between approximately 8 and 25 $\mu$eV. Horizontal dashed lines represent phonon mode energies at 2, 4, and 6 GHz, and the dash-dotted red line marks a typical superconducting resonator frequency at 5.2 GHz.
@@ III Phonon-Assisted $T_{1}$ Relaxation
Figure 4: Phonon-assisted relaxation process in a Ge qubit. An excited spin state relaxes to the ground state via spontaneous phonon emission with energy matching the Zeeman splitting $\Delta E$. This process is enabled by spin-orbit coupling and influenced by quantum dot geometry, crystal orientation, and phonon mode structure.
@@
Figure 5: Comparison of phonon bottleneck and cavity-enhanced relaxation regimes. Left: In small quantum dots, relaxation is suppressed when phonon wavelengths exceed the dot size. Right: A PnCC enhances spin-phonon coupling by concentrating phonon density at the qubit’s transition energy.
@@
[…]
@@ IV Phononic Crystal Cavity for Control and Coupling
Figure 6: Schematic of a Ge quantum dot coupled to a PnCC. The cavity confines GHz-frequency acoustic phonon modes that can be electrically excited to interact with qubit states. Electric field control enables spectral tuning of the spin splitting into resonance with the phonon mode.
@@
where $\sigma_{+}^{(i)}=|1\rangle_{i}\langle 0|$ and $\sigma_{-}^{(i)}=|0\rangle_{i}\langle 1|$ are the spin raising and lowering operators for qubit $i$.
@@
To support this form of coupling in Ge-based hole-spin qubits, we have designed a two-dimensional phononic crystal with an absolute bandgap centered around 4–6 GHz, with a bandwidth of 2–3 GHz. This frequency range corresponds to thermal energies near 0.3 K and is well-matched to dilution refrigerator environments operating at ∼$\sim 100\,\text{mK}$. At these temperatures, thermal occupation of acoustic modes can lead to decoherence. The PnCC structure acts as a spectral filter, suppressing thermally populated phonon modes outside the desired range while confining coherent phonon modes for interaction. This selective filtering enhances coherence and supports long-range, phonon-mediated qubit control in scalable Ge quantum architectures.
@@
Figure 9: Spatial overlap profile between the phononic displacement field at 6 GHz and a Gaussian-modeled hole-spin qubit envelope. The significant overlap highlights spatial compatibility essential for maximizing $g_{sp}$.
@@
In this work, we estimate the relative strength of spin–phonon coupling $g_{\text{sp}}$ by evaluating the spatial overlap between the phononic strain field and the qubit’s envelope wavefunction. To model the spatial envelope of the hole-spin qubit, we assume a Gaussian wavefunction of the form
@@
Target Q-Factor $>10^{4}$; long phonon lifetime
@@
Overlap Integral ∼$\sim 2.6\times 10^{-2}$; at 6 GHz
@@
where $v_{s}\approx 5000\,\mathrm{m/s}$ is the longitudinal sound velocity in Ge. For instance, targeting $f_{\text{center}}=5\,\mathrm{GHz}$ yields $a\approx 1.0\,\mu\mathrm{m}$. The hole radius should be selected in the range $r=(0.2\text{–}0.4)\,a$ to achieve a wide and robust acoustic bandgap.
@@ V Experimental Considerations
To mitigate hyperfine-induced decoherence, isotopically enriched 74Ge substrates are used to eliminate nuclear spins from the lattice [9]. This isotopic purification enables observation of long spin coherence times, often exceeding 100 $\mu$s, as demonstrated in Ramsey and Hahn-echo experiments [8, 9].

### v3

@@ Abstract
We present a design and modeling of a scalable quantum processor architecture utilizing hole-spin qubits defined in gate-controlled germanium (Ge) quantum dots, where coherent spin-phonon coupling is predicted to facilitate qubit manipulation and long-range interactions. The architecture exploits the strong, electrically tunable spin-orbit interactions intrinsic to hole states in Ge, integrated with high-quality phononic crystal cavities (PnCCs) to enable fully electrical qubit control and phonon-mediated coupling. Employing a streamlined simulation framework built upon multiband $\mathbf{k}\cdot\mathbf{p}$ modeling and finite-element methods, we quantify key performance metrics, including electrically tunable $g$-factors ranging from $1.3$ to $2.0$, spin-phonon coupling strengths up to $6.3\,\mathrm{MHz}$, phononic cavity quality factors exceeding $10^{4}$, and phonon-mediated spin relaxation times ($T_{1}$) reaching milliseconds. The proposed architecture concurrently achieves extended spin coherence and rapid gate operations through strategic electric field modulation and engineered phononic bandgap environments. Furthermore, isotopically enriched, high-purity Ge crystals significantly enhance device coherence by minimizing disorder and hyperfine interactions. This integrated approach, merging advanced materials engineering, precise spin-orbit coupling, and phononic cavity design, establishes a promising CMOS-compatible pathway toward scalable, high-fidelity quantum computing.
@@ I Introduction
Figure 1: Comparison of representative qubit technologies based on coherence time and gate speed, reflecting current experimental capabilities [1, 2, 8, 9]. Ge-based hole-spin qubits with phonon coupling are positioned in a favorable regime that combines fast, all-electrical control with long coherence, underscoring their promise for high-fidelity quantum information processing. Ge-phonon-coupled reflects two-qubit rates mediated by a phonon bus (dispersive exchange), typically $\sim 0.5\text{--}1~\mu\mathrm{s}$, whereas single-qubit in Ge remains $\mathcal{O}(10\text{--}50\,\mathrm{ns})$ under our operating conditions (see Sec. IV).
Within this landscape, we benchmark Ge hole–spin qubits against leading alternatives beyond superconducting circuits and trapped ions, namely neutral atoms in optical tweezers (including Rydberg–mediated entangling gates) and solid–state defect centers (e.g., NV centers in diamond). These modalities achieve competitive coherence and high–fidelity control but impose distinct hardware and integration constraints (optical access, cryogenic operation, and packaging), which set their operating points in the gate–speed/coherence plane. To make the comparison concrete, Fig. 1 includes representative single– and two–qubit time scales and coherence values for these platforms; the caption reports typical experimental ranges to provide context for the Ge results highlighted here.
@@
A critical factor influencing the performance of Ge-based hole-spin qubits is material quality, particularly impurity scattering and charge noise, which significantly degrade qubit coherence. To address these challenges, this study leverages in-house capabilities at the University of South Dakota (USD) for growing detector-grade, high-purity Ge crystals, achieving impurity concentrations as low as $\sim 10^{10}\,\text{cm}^{-3}$ [13, 14, 15, 16, 17, 18, 19]. This vertically integrated crystal growth infrastructure provides a robust pathway for fabricating low-disorder quantum dots, thus enabling detailed exploration of phonon-assisted spin manipulation within exceptionally clean Ge-based systems.
A distinctive strength of our proposed architecture lies in the strategic integration of isotopically enriched, high-purity 74Ge substrates, the strong and electrically tunable spin-orbit coupling inherent to hole states, and precisely engineered phononic crystals (PnCs) with embedded PnCCs. Together, these components form a robust platform for spin-phonon-based quantum information processing. The PnC structure is designed to surround the quantum dot, creating a phonon bandgap that suppresses unwanted phonon modes and isolates a target frequency range (e.g., around 6 GHz). Localized defects introduced into the PnC form PnCCs that support discrete acoustic modes resonant with the desired frequency. These modes are spatially aligned with the wavefunction of the hole-spin qubit in the Ge quantum dot, enabling coherent spin-phonon interaction via the deformation potential. Related theory for Ge spins embedded in quasi-2D phononic crystals shows that tuning the Zeeman splitting into a phononic bandgap suppresses one-phonon relaxation and enables long-range coupling via virtual phonon exchange, which we extend here to Ge hole-spin dots [20]. While the PnC acts as an acoustic shield to minimize phonon leakage and external noise, the PnCC enhances the local phonon density of states, thereby increasing the coupling strength. This integrated approach facilitates strong, localized spin-phonon coupling essential for coherent qubit control, high-fidelity readout, and phonon-mediated qubit entanglement.
@@ II Spin-Phonon Coupling in Germanium Quantum Dots
Figure 2: Schematic showing spin-phonon coupling in a Ge quantum dot. Left: In zero magnetic field, the hole-spin states are degenerate. Middle: With an external magnetic field applied along [001], the states split by the Zeeman energy $\Delta E=g\mu_{B}B$; phonons with energy $E_{\rm ph}=\hbar\omega=\Delta E$ can induce spin-flip transitions via spin–orbit coupling, which are enhanced by heavy-hole/light-hole mixing in low-symmetry environments. Here, $\lvert v\rangle$ and $\lvert w\rangle$ denote the spin-down and spin-up eigenstates of the spin qubit in the Ge quantum dot. Right: A hole initially in $\lvert v\rangle$ (spin-down) can absorb a matching-energy phonon and transition to $\lvert w\rangle$ (spin-up).
@@
At cryogenic temperatures ($1~\text{K}$) relevant for quantum computing, acoustic phonons dominate over optical modes. The primary spin relaxation mechanism in this regime is acoustic deformation potential coupling, whereby phonons modulate the local band structure and mediate transitions between spin states [7, 9]. This spin-phonon interaction plays a dual role: while it enables efficient spin initialization and readout via phonon emission or absorption, it can also introduce decoherence through relaxation ($T_{1}$) and dephasing ($T_{2}^{*}$) pathways.
[…]
@@ III Phonon-Assisted $T_{1}$ Relaxation
Figure 4: Phonon-assisted relaxation process in a Ge qubit. Left: In an external magnetic field, the excited spin state relaxes to the ground state via spontaneous emission of a phonon with energy $E_{\rm ph}=\hbar\omega=\Delta E$ (center). This one-phonon process is enabled by spin–orbit coupling and is influenced by heavy-hole/light-hole mixing as well as device-specific factors—quantum-dot geometry, crystal orientation, and the phonon-mode structure (bulk substrate, phononic crystal, or cavity).
@@
Figure 5: Comparison of phonon bottleneck and cavity-enhanced relaxation regimes. Left: In small quantum dots, when the dominant phonon wavelength $\lambda_{\rm ph}$ exceeds the dot size $L$, the spin–phonon matrix element is suppressed by the dot form factor (the “phonon bottleneck”), yielding longer $T_{1}$. This panel is the reference (no phononic crystal). Right: A 2D phononic-crystal cavity (PnCC) creates a localized defect mode inside an engineered bandgap, concentrating the phonon density of states at the qubit transition $(\hbar\omega_{Z}=\Delta E)$ and thereby enhancing spin–phonon coupling; off resonance (within the bandgap) relaxation remains suppressed, while on resonance with the defect mode it can be intentionally increased or used for engineered interactions. The diagonally hatched gray band marks the phononic bandgap where propagating acoustic modes are suppressed, reducing $T_{1}^{-1}$. The color scale and axes are identical in both panels for direct comparison.
@@
Under optimized gating and environmental control, we estimate that $T_{2}^{*}$ in this architecture could exceed $1\text{--}10~\mu\text{s}$, consistent with or exceeding values reported in other planar Ge and Si-based platforms [8, 31]. Such coherence times are compatible with the simulated spin relaxation time $T_{1}\sim\text{ms}$, and sufficient to support high-fidelity spin readout, control, and phonon-mediated entanglement operations.
Beyond the inhomogeneous dephasing time, the intrinsic coherence time $T_{2}$ provides a more fundamental measure of phase coherence, representing the timescale for irreversible decoherence processes such as spin-bath interactions or high-frequency phonon scattering. $T_{2}$ is typically extracted using spin-echo or dynamical decoupling protocols that refocus slowly varying noise components [5, 34]. In our proposed design, several features are expected to positively influence $T_{2}$: the use of isotopically enriched Ge eliminates nuclear-spin-induced dephasing; the phononic bandgap suppresses incoherent phonon scattering; and cavity-induced spectral narrowing reduces residual environmental coupling [36, 35]. Although detailed modeling of $T_{2}$ is beyond the present scope, we anticipate that echo-based coherence times in the range of $50\text{--}100~\mu\text{s}$ are achievable, consistent with recent experimental results [31, 37, 38]. Furthermore, the incorporation of charge-noise-insensitive “sweet spots” [30] and advanced decoupling sequences may further extend $T_{2}$, positioning this platform for high-fidelity, scalable quantum operations mediated by phonons.
@@ IV Phononic Crystal Cavity for Control and Coupling
Figure 6: Phononic-crystal cavity concept and qubit coupling. Left: Gate-defined Ge hole-spin quantum dot with adjacent control/readout gates; the dot is positioned near a displacement antinode to maximize the spin–phonon coupling $g_{\rm sp}$. Middle: A point defect in a 2D phononic crystal confines a GHz mechanical mode inside the bandgap; the localized strain field overlaps the dot and can be driven by gate modulation, enabling interaction with the spin via spin–orbit coupling. Right: DC electric-field control ($E_{z}$) tunes the effective $g$-factor and Zeeman splitting $\Delta E=\hbar\omega_{Z}$, allowing spectral alignment (or deliberate detuning) with the cavity mode for controlled coupling and relaxation.
@@
where $\sigma_{+}^{(i)}=|1\rangle_{i}\langle 0|$ and $\sigma_{-}^{(i)}=|0\rangle_{i}\langle 1|$ are the spin raising and lowering operators for qubit $i$. Eq. 3 follows from the deformation-potential (Bir–Pikus) spin–phonon Hamiltonian after quantizing the elastic field and projecting onto the qubit subspace; see Refs. [43, 44, 45, 46] for detailed derivations in related solid-state platforms.
@@
To support this form of coupling in Ge-based hole-spin qubits, we have designed a two-dimensional phononic crystal with an absolute bandgap centered around 4–6 GHz, with a bandwidth of 2–3 GHz. This frequency range corresponds to thermal energies near 0.3 K and is well-matched to dilution refrigerator environments operating at $\sim 100\,\text{mK}$. At these temperatures, thermal occupation of acoustic modes can lead to decoherence. The PnCC structure acts as a spectral filter, suppressing thermally populated phonon modes outside the desired range while confining coherent phonon modes for interaction. This selective filtering enhances coherence and supports long-range, phonon-mediated qubit control in scalable Ge quantum architectures.
@@
Figure 9: Spatial overlap profile between the phononic displacement field at 6 GHz and a Gaussian-modeled hole-spin qubit envelope. Normalized overlap amplitude $\mathcal{O}_{\rm norm}=\left|\int\varepsilon(\mathbf{r})|\psi(\mathbf{r})|^{2}d^{3}r\right|\Big{/}\Big{[}\Big{(}\int|\varepsilon(\mathbf{r})|^{2}d^{3}r\Big{)}^{1/2}\!\Big{(}\int|\psi(\mathbf{r})|^{4}d^{3}r\Big{)}^{1/2}\Big{]}$. Larger values indicate stronger spin–phonon coupling for fixed material tensors. The significant overlap highlights spatial compatibility essential for maximizing $g_{sp}$.
@@
In the acoustic regime relevant here, the Bir–Pikus perturbation renders $H_{\rm so}[u(\mathbf{r})]\propto\varepsilon(\mathbf{r})$, so the matrix element $M_{\rm sp}$ is dominated by the spatial overlap $\int d^{3}r\,\varepsilon(\mathbf{r})\,|\psi(\mathbf{r})|^{2}$ (up to known material tensors), which motivates the ‘overlap’ metric used below to compare different PnC designs. To model the spatial envelope of the hole-spin qubit, we assume a Gaussian wavefunction of the form
@@
Target Q-Factor $10^{4}$; long phonon lifetime
@@
Overlap Integral $\sim 2.6\times 10^{-2}$; at 6 GHz
@@
where $v_{s}\approx 5000\,\mathrm{m/s}$ is the longitudinal sound velocity in Ge. For instance, targeting $f_{\text{center}}=5\,\mathrm{GHz}$ yields $a\approx 1.0\,\mu\mathrm{m}$. The hole radius should be selected in the range r=(0.2r=(0.20.4)a0.4)\,a to achieve a wide and robust acoustic bandgap, consistent with prior bandgap maps of 2D phononic membranes where complete gaps maximize near filling fractions corresponding to $r/a\approx 0.25$$0.40$ for triangular/honeycomb lattices [50, 51, 52, 53, 20].
@@
In practice, this filtering–and–control architecture yields three concrete benefits that reinforce the goals above. First, because spin–phonon coupling in Ge is typically stronger than spin–photon coupling, on-chip acoustic actuation (IDTs/SAWs) enables efficient, local qubit control while reducing external hardware. Second, engineered phononic bandgaps act as spectral filters that suppress off-resonant environmental phonons, protecting both $T_{1}$ and $T_{2}^{\ast}$ while transmitting the targeted resonant mode for coherent interaction. Third, a guided phonon bus provides long-range, reconfigurable connectivity between distant dots without tight lithographic proximity constraints, supporting scalable architectures for high-fidelity operations.
@@ V Experimental Considerations
To further minimize lattice mismatch and optimize epitaxial growth quality, we employ a relaxed Si1-xGex virtual substrate with Ge content $x\approx 0.95\text{--}1.0$. This composition closely matches the lattice constant of pure Ge, thereby suppressing strain-induced dislocations and supporting the formation of atomically smooth, low-defect Ge quantum wells. The reduced lattice mismatch not only improves crystal quality but also enhances hole mobility and uniformity of confinement potential, both of which are critical for high-coherence qubit operation and low-noise detector performance. This materials platform provides an ideal foundation for the growth of isotopically enriched 74Ge layers and the fabrication of low-disorder quantum dot and phononic crystal structures essential for scalable quantum architectures.
@@
Quantitatively, the thermal occupation that drives these trends is $n_{B}(\omega,T)=\big{[}\exp(\hbar\omega/k_{B}T)-1\big{]}^{-1}.$ For a 6 GHz mode, $n_{B}(100\,\mathrm{mK})\approx 0.06$ and $n_{B}(300\,\mathrm{mK})\approx 0.6$. Consistent with the discussion above, $g_{sp}$ is essentially temperature-independent, but the increased $n_{B}$ at higher $T$ opens additional stimulated emission/absorption channels that shorten $T_{1}$ and load the cavity, reducing its effective $Q$ even though the bandgap itself remains. Accordingly, we target $T\lesssim 100\,\mathrm{mK}$ for phonon-assisted readout and coupling; operation at $200\text{--}300\,\mathrm{mK}$ is still feasible for some tasks but with explicitly reduced $Q$ and elevated $n_{B}$, as now stated in the Methods.